OpenStudy (anonymous):

Calculate the producers' surplus for the supply equation at the indicated unit price p. (Round your answer to the nearest cent.) p = 7 + 2q^1/3; p = 18

OpenStudy (dumbcow):

producer surplus = area above curve and below given price \[= PQ - \int\limits_0^Q (7+2q^{1/3}) dq\] P = 18 Q = 1331/8

OpenStudy (dumbcow):

after integrating \[\rightarrow (18)(\frac{1331}{8}) - (7q +\frac{3}{2}q^{4/3}) |_0^{1331/8}\] \[= \frac{9}{4}(1331)- \frac{7}{8}(1331)-(\frac{3}{16})(\frac{11}{2})(1331)\] \[= \frac{11}{32}(1331)\] = 457.53